OpenStudy (anonymous):

Please explain to me how to multiply and simplify: x^2-4/x^2-1 times x+1/x+2

OpenStudy (anonymous):

x^2-4= (x+2)(x-2) x^2-1=(x+1)(x-1) there fore (x+2)(x-2) x+1 ----------- * ------- (x+1)(x-1) x+2 x-2 ------ x-1

OpenStudy (anonymous):

thanks you sooo much for explaing (:

OpenStudy (anonymous):

ur welcome sir